Review Part A Objects A and B are both positively charged. Both have a mass of 250 g, but A has twice the charge of B. When A and B are placed 60 cm apart, B experiences an electric force of 0.50 N What is the charge on A? Express your answer to two significant figures and include the appropriate units. qA- 1 Value Units Submit|F Request Answer Part B If the objects are released, what is the initial acceleration of A? Express your answer to two significant figures and include the appropriate units. a= 1 Value Units
Answer #1

Apply coulomb's law

F= kqAqB/r^2

as per given data

let charge on B is Q then charge on A is 2Q

F = k ( 2Q) (Q)/r^2

= k 2Q^2/r^2

Q = sqrtFr^2/2k

= sqrt (0.5)((0.6)^2/2 * 9* 10^9

=3.16* 10^-6 C

charge on A is 2Q = 6.3* 10^-6 C

(b)

Apply Newton second law

F = ma

a = F/m = 0.5/0.250 kg = 2 m/s^2